A computer password must be eight characters long. How many passwords are possible if only the 26 letters of the alphabet are allowed?

explain with details
Answer #1

solution :

Password is of eight characters and for each eight character, there is possibility of 26 letters hence number of possible password are 26×26×26×26×26×26×26×26 = 268

if repeatation is not allowed

Then the number of passwords possible = 26*25*24*23*22*21*20*19 = 62990928000